OpenStudy (anonymous):

how do I find the zeros of x^4 - x^3 -2x^2 algebraically?

OpenStudy (anonymous):

well that's what got

OpenStudy (e.mccormick):

You can factor it or use assorted tests to find possible ones and test them with division. Those are two possible ways.

OpenStudy (anonymous):

my book says x= -1, 0, 0, 2

OpenStudy (e.mccormick):

The 0 should be obvious because all terms have x in them.

OpenStudy (e.mccormick):

\( x^4 - x^3 -2x^2 \implies \) \(x^2( x^2 - x -2) \) See, the \(x^2\) out front makes for the 0 parts of your answer. Then you just need to factor the part in the ( ) to get the others. When something says to find the roots, factors, zeros, x intercepts, or solutions for = 0, they are all asking for the same thing.

OpenStudy (anonymous):

thankyou

OpenStudy (e.mccormick):

Do you understand the rest of how to do these?

OpenStudy (e.mccormick):

This is an example and NOT your problem: Lets say I had the factored form of \((x-5)(x+3)x\). Then the roots/zeros are when that = 0. \((x-5)(x+3)x=0\) means: \(x-5=0\) or \(x+3=0\) or \(x=0\) So by solving each of those factors for x, I know the roots/zeros. \(x-5+5=0+5\) or \(x+3-3=0-3\) or \(x=0\) \(x=5\) or \(x=-3\) or \(x=0\) So the roots/zeros are at \(x=-3,0,5\) Well, what if I was given: \(x^3-2x^2-15x\) It would be the same. See, that is just the non-factored form of the same thing. That is why factoring these really helps. That is why I started by factoring yours.